the length of a rectangle is 20ft. less than twice the width. of the perimeter of the rectangle is 260ft. find the length and width of the rectangle

Answer by mathstutor494(120)

Let width = x
Then length = 2x-20
Perimeter of the rectangle = 2(Length+Width)
2( 2x-20+x) = 260
2(3x-20) = 260
3x-20= 130
3x = 130+20
x = 150/3 => 50
Therefore Length = 2x-20 => 2*50-20 => 100-20 => 80
Ans : Length 80 ft, Width 50 ft